OpenStudy (anonymous):

What would the derivative of the area of a right triangle formula be? A=(1/2)b*h

OpenStudy (mathstudent55):

With respect ot which variable?

OpenStudy (anonymous):

With respect to time sorry!

OpenStudy (mathstudent55):

Do both the base and height vary with respect to time?

OpenStudy (anonymous):

Yes, the problem is A 13 ft ladder is leaning against a house when its base begins to slide away. By the time the base is 3 ft from the house, the base is moving at the rate of 2 ft/sec. Answer the following: and I need to find: At what rate is the area of the triangle formed by the ladder, wall, and ground changing then? But I think I messed up differentiating the area of the triangle!

OpenStudy (mathstudent55):

y = uv y' = uv' + vu'

OpenStudy (anonymous):

Yes. So is A'=(1/2)b'*h + (1/2)b*h' correct?

OpenStudy (anonymous):

Pretending the ( )' = d( )/dt

OpenStudy (mathstudent55):

\(A = \dfrac{1}{2}bh\) \(A = \dfrac{1}{2}(bh)\) \(\dfrac{dA}{dt} = \dfrac{1}{2}\dfrac{d}{dt}(bh)\) \(\dfrac{dA}{dt} = \dfrac{1}{2} \left(b \dfrac{dh}{dt} + h \dfrac{b}{dt} \right)\)

OpenStudy (mathstudent55):

Yes, you are correct.

OpenStudy (mathstudent55):

Sorry, last fraction above in last line should be db/dt

OpenStudy (anonymous):

Okay awesome thanks! :-)

OpenStudy (mathstudent55):

yw